OpenStudy (unklerhaukus):

Show that: \[\cos^2 (nx) = \frac{1 + \cos (2nx)}2\]

OpenStudy (astrophysics):

Does 1+cos(2x) = 2cos^2(x) not work here?

ganeshie8 (ganeshie8):

I would start with \(\cos^2nx+\sin^2nx=1\)

OpenStudy (unklerhaukus):

\[LHS=1-\sin^2nx\]

OpenStudy (unklerhaukus):

@Astrophysics how can i show: \[1+\cos(2x) = 2\cos^2(x)\]

OpenStudy (amilapsn):

@UnkleRhaukus use the identity for \(\cos(A+B)\)

ganeshie8 (ganeshie8):

\[1+\cos(2x) = \mathcal{R}(1+e^{i(2x)}) = \mathcal{R}(1+e^{ix}e^{ix}) \\~\\= \mathcal{R}(1+(\cos x+i\sin x)(\cos x+i\sin x))\\~\\ =\cdots \]

OpenStudy (astrophysics):

You can use cos(A+B) = cosAsinB+sinAcosB cos(2x) = cos(x+x) = .... you see :P

OpenStudy (unklerhaukus):

\(\mathcal R\) is \(\Re\), the real component ?

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):